# Formal Analysis of the Contract Automata Runtime Environment with Uppaal: Modelling, Verification and Testing

### v1

@@ 5. Formal model
The formal model of CARE is described. All models used in this paper together with the evaluated formulas are available in [6]. The network of automata is composed of a RunnableOrchestration automaton modeling the orchestrator, and each service is modeled by two automata: RunnableOrchestratedContract and SocketTimeout. Figure 5 shows the system declarations of the model. In particular, there is a single instance of RunnableOrchestration whilst there are N instance of RunnableOrchestratedContract and SocketTimeout. Indeed, both RunnableOrchestratedContract and SocketTimeout templates take as parameter id_t, which is a special global declaration of Uppaal (namely, typedef int[0,N-1] id_t) to instruct the model to produce N replicas of these template, where as stated above N is also a global constant of the model.
@@
Figure 6 displays the template automaton for the RunnableOrchestration (i.e., the orchestrator), while the template automaton for the RunnableOrchestratedContract (i.e., the service) has as parameter the id of the service (of type id_t) and is depicted in Figure 7. The behaviour according to the given configuration of action and choice is modelled inside each automaton. We anticipate that we will use variants of these two automata in Section 6, in order to perform different analyses. In particular, in Figures 6 and 7 the configurations of choice and action (for both the services and the orchestrator) are instantiated non-deterministically. Another version of the model is also available where the configuration options are also parameters of the templates. The repository [6] contains the model equipped with traceability information, i.e., each transition of the model has a comment describing the class and the lines of the source code that correspond to the specific behaviour of the transition. The models used for generating tests, together with the generated tests, are also available in [6].
@@ 5.1. Description of the automata
We now briefly describe Figure 6 and Figure 7. The initial state is depicted with a double circle. In the first transition of the orchestrator, one of four options encoding the combinations of choice and action is selected non-deterministically. The function initialize(conf) instantiates accordingly the choice and action global variables such that all have the same configuration, and also initialises the communication buffers.
From the location CheckCompatibility of the orchestrator a loop is executed to communicate with all services to check if all have the same configuration. The orchestrator sends the message ORC_CHECK and its configuration (action and choice). If a service has the same configuration an ACK is sent, or an ERROR message otherwise. In case of no errors, the orchestration starts. From the location Start the orchestrator internally decides (based on the orchestration) whether to perform a choice or an action. The choice of termination is modelled as a third alternative. As stated previously, after a choice is completed the orchestrator moves to a state where only an action or termination can be chosen. After an action is completed, the orchestrator returns to the Start location.
@@
In the case of termination, the orchestrator sends to all services an ORC_STOP message and the orchestration terminates. In the case of a choice, firstly all services receive the message ORC_CHOICE. If the choice is dictatorial, the orchestrator decides autonomously, and no further interactions are necessary. Otherwise, in the case of a majoritarian choice, the services involved in the choice are selected non-deterministically, and a message ORC_CHOICE is sent to them. Concretely, the involved services are those who perform an action in one of the outgoing transitions from the current state of the orchestration. The services that are not involved will receive a SKIP message. The involved services then receive from the orchestrator the available choices (i.e., the forward star of the current state). These concrete choices are abstracted by the constant CHOICES. Each involved service now replies with its choice, abstracted as a message SERVICE_CHOICE. After all involved services have voted, the orchestrator will decide accordingly.
@@ 6. Analysis
We now describe the analyses we performed on the model. The modelling activity led to some issues in both the implementation and the model, which were all fixed (we remark that an already existing application was modelled). Other formal checks on the model were performed to ensure the properties described in this section (e.g., absence of deadlocks, absence of orphan messages). We have used Uppaal version 4.1.26-1, February 2022.
@@ 6.1. Validation through modelling
On a side note, thorough testing is generally more time-consuming than designing a formal model similar to the one in this paper, and Uppaal has been used to automatically generate tests from the formal model.
@@ 6.2. Formal Verification
Performances. Statistical model checking has been used to scale to larger systems, and the verification has been performed in a few seconds on a standard laptop. Conversely, exhaustive model checking necessitated more resources and has been limited to smaller configurations generating hundreds of millions of states (see below). In this case, the verification has been performed on a machine with Intel(R) Core(TM) i9-9900K CPU @ 3.60 GHz equipped with 32 GB of RAM. Uppaal has been configured for maximising the state space optimization and reusing the generated state space. Logs of the experiments are available in [6]. The summaries of the analysis performed with statistical (resp., standard) are in Table 1 (resp., Table 2).
@@
Table 2. A summary of the properties analysed with standard model checking that are satisfied by the model
@@ 6.3. Parameters Tuning
The verification process involves employing specific parameter configurations within the model. This encompasses various aspects, such as setting the delays in reading and writing, setting socket timeout thresholds, adjusting buffer sizes, assigning probability weights, and determining the number of services involved (i.e., the instantiations of the RunnableOrchestratedContract template). For deriving the desired set-up of parameters, we employ statistical model checking. If not stated otherwise, the parameters $\alpha$ and $\varepsilon$ of the statistical model checker are set to $0.05$ and $0.005$, respectively (see Section 3).
@@
It is crucial to note that we do not employ statistical model checking to determine values (such as buffers size) for use in the concrete implementation. Instead, these quantified values are used solely within the model. For instance, in the actual implementation, the size of Java TCP/IP socket buffers is fixed (more below). Our objective is to employ parameter configurations that ensure realistic modelling and improve the performances of the exhaustive model checking. Realistic modeling entails accurately representing the behavior of the real system, by reducing the probability of filling the buffers, timeouts, or excessive communication delays. Failure to maintain these conditions could potentially invalidate the results of the formal verification. Improving performances, on the other hand, entails reducing the state space of the model.
Firstly, the values of the probabilities weights pstop, pchoice, pnochoice, and paction (see Figure 6) can be tuned based on average values extracted from the orchestrations subject of the analysis. Indeed, as stated in Section 4, the underlying orchestration automaton is abstracted away. Note that, e.g., in location Start the probability of performing a choice is $\frac{\tt pchoice}{\tt pchoice+pnochoice}$ (pchoice is the probability of performing a choice and pnochoice is the probability of not performing a choice). For example, the orchestration in Figure 8 (right) can be modelled by tuning the probabilities to pstop=25, pchoice=1, pnochoice=0, and paction=75.
Figure 8. On the left side, the SocketTimeout automaton. On the right side, the orchestration automaton taken from the composition service example in [11]
Next, we address the buffer size (denoted as variable queueSize). It is important to note that the buffers in the model are represented by global arrays utilized by the automata for enqueuing and dequeuing values. These global arrays serve as models of the actual buffers found in Java TCP/IP sockets, where the default size is typically 8 KB. Our objective is to prevent unnecessary growth in the model’s state space while ensuring a low probability of the buffers filling up, similar to the behavior observed in the real application. The formula:
@@
Next, we consider the real-time behavior of the model and focus on determining appropriate values for the rates write and read, which represent the delays in writing to and reading from a buffer, respectively. The average message delay in Java TCP/IP sockets is affected by multiple factors, such as network conditions and server load. Delays can be sampled using tools like tcpdump, and the rate can be estimated by calculating the inverse of the sample mean. Additionally, we consider the variable timeout, which represents the timeout threshold. Our aim is to achieve three objectives. First, we strive to maintain a low probability of encountering timeouts. Second, we seek to ensure a high probability of terminating within a specific timeframe, which we have set to be 500 time units, corresponding to the duration used in our experimental setup. Third, we aim to keep the timeout threshold at a lower value in order to decrease the state space and facilitate model checking.
@@
Concerning the instances of the templates, in all experiments there is one instance of the orchestrator template and either 4 or 5 instances of the service template. For the exhaustive model checking phase, we used two small configurations of (number of services, buffers size). The first is c1=(4,5), the second is c2=(5,3). In fact, the configuration (5,4) remained inconclusive in the experiments due to the need for generating billions of states and the inadequate memory capacity of the utilized machine. The verification of larger configurations necessitates either relying exclusively on statistical model checking or employing more powerful machines.
@@ 6.4. Verification
Once the model’s parameter configuration is determined, our next step involves verifying additional formal properties.
@@
i.e., if the orchestrator decides to terminate then eventually all services and the orchestrator terminate. The formula p-->q is a shortcut for A[](p imply A<>q). Thus, this formula states that for all executions and for all states either the orchestrator is not in the location ror.Stop or all executions passing through that location will eventually lead to a state where all services and the orchestrator have terminated or a timeout failure is experienced. The formula holds in both configurations c1 and c2. The first (resp. second) configuration required roughly 3 (resp. 30) minutes, 1.5 (resp. 12) Giga of memory, and explored 52 (resp. 426) million states.
@@
the formula states that for all executions and for all states of the composed system, either there is always at least one enabled transition or either a timeout failure has been experienced or all services and the orchestrator are in the Terminated location. In the formula, not deadlock is a special predicate provided by Uppaal. As expected, this property is satisfied in both configurations. The first (resp. second) configuration required roughly 3 (resp. 40) minutes, 1.5 (resp. 13) Giga of memory, and explored 25 (resp. 189) million states. This also proves that in a correct configuration the Error location is never reached, because this would result in a deadlock and the above property would not be satisfied. This property is also satisfied when the size of the buffers is $3$. However, if we further reduce the size, then a deadlock occurs. This is because from location CheckCompatibility the orchestrator requires to insert three messages in the buffer of the receiver in one step. By dividing these three send operations in three non-committed transitions it is possible to further reduce the buffer size.
@@
this property quantifies the probability of termination with at least one message remaining in any buffer. To verify whether all buffers are empty, we utilize the predicate allEmpty(). As expected, the probability is found to be nearly zero. We proceed with an exhaustive verification by employing the property
@@
the above formula can be read as follows: in all states of all executions either all buffers are empty or someone has not terminated yet. The above property is valid in both configurations, as expected. The first (resp. second) configuration required roughly 2 (resp. 27) minutes, 1.5 (resp. 12.5) Giga of memory, and explored 25 (resp. 189) million states. It is also possible to verify that there is no dummy execution in which the services and the orchestrator never interact (and thus all buffers are trivially empty). This can be verified with the formula:
@@
that, as stated above, checks if there exists an execution where all states have empty buffers (excluding the dummy execution scenario in which the timeout occurs at the beginning). As expected, this property is not valid in the model, for both configurations. The first (resp. second) configuration required roughly 3 (resp. 24) seconds, 1.4 (resp. 10) Giga of memory, and explored 24 (resp. 28) states.
No interference. When discussing the distributed match action in Section 5, we stated that it is never the case that some service interferes in a match in which it is not involved. This guarantees that it is safe to use two one-position buffers for all communications between any two services involved in a match. To verify this, we perform statistical model checking of the following formula
@@ 7. Testing
The abstract test cases generated by Uppaal have been transformed into concrete JUnit tests for the source code. Uppaal provides various methods for test generation (see Section 3). We utilized the generation from reachability queries (i.e., of the form E<>). In our models, these queries encode specific simulation traces that are relevant to the specific orchestration employed in the tests and are designed to cover all transitions of the model. The specific simulation traces also guide the generation of the concrete test cases from the abstract test cases. To achieve this, two global variables, namely steps[] and step, are introduced in the models to encode the desired orchestration in the query. The length of the array steps[] is equal to the length of the trace being simulated, and the variable step is a pointer to the current position in the array. Whenever a specific transition is executed, the variable steps[step] is updated with a value encoding the transition being executed, and the counter step is incremented. For example (see Figure 9), whenever the transition reaching state Error is executed, the value ERROR is assigned to steps[step]. As another example, whenever a centralised match transition is executed by the orchestration, the value CENTRALISED_MATCH is assigned to steps[step], encoding the execution of the request action in the match (see Section 5). In Section 7.3, we will provide an example showing how these additional variables are used in a query.

### v2

@@ 5. Formal model
This section describes the formal model of CARE. All models used in this paper together with the evaluated formulas are available in [10]. The network of automata is composed of a RunnableOrchestration automaton modelling the orchestrator, and each service is modeled by two automata: RunnableOrchestratedContract and SocketTimeout. Figure 7 shows the system declarations of the model. In particular, there is a single instance of RunnableOrchestration whilst there are N instance of RunnableOrchestratedContract and SocketTimeout. Indeed, both RunnableOrchestratedContract and SocketTimeout templates take as parameter id_t, which is a special global declaration of Uppaal (namely, typedef int[0,N-1] id_t) to instruct the model to produce N replicas of these template, where as stated above N is also a global constant of the model.
@@
Figure 8. A simplified schema of the RunnableOrchestration automaton
Figure 9 displays the template automaton for the RunnableOrchestration (i.e., the orchestrator), while the template automaton for the RunnableOrchestratedContract (i.e., the service) has as parameter the id of the service (of type id_t) and is depicted in Figure 10. Recall that, for readability, the Timeout location is not displayed (see Section 4.1). The behaviour according to the given configuration of action and choice is modelled inside each automaton. We anticipate that we will use variants of these two automata in Section 6, in order to perform different analyses. In particular, in Figure 9 the configurations of choice and action (for both the services and the orchestrator) are instantiated non-deterministically by the transition outgoing the initial state (see Section 5.1). Note that, in this case, the variables choice and action are not parameters of the templates. Another version of the model is also available, where the configuration options, instead of being selected non-deterministically, are fixed as parameters of the templates. Note that in Uppaal templates cannot be primitively instantiated non-deterministically. The repository [10] contains the model equipped with traceability information, i.e., each transition of the model has a comment describing the class and the lines of the source code that correspond to the specific behaviour of the transition. The models used for generating tests, together with the generated tests, are also available in [10].
@@ 5.1. Description of the automata
Before discussing the two automata, we describe a simplified template, depicted in Figure 8.
5.1.1. Simplified template
Note that Figure 8 is not a Uppaal automaton. The initial state is depicted with a double circle. The first activity consists in checking whether all services and the orchestrator have the same configuration (CheckCompability). If this check is successful, the orchestration starts. From the location Start, the orchestrator internally decides (based on the orchestration) whether to perform a choice or an action. The choice of termination is modelled as a third alternative. Probability weights are used to model the probability of performing a choice, an action or to terminate (pchoice, pstop, paction). As stated previously, after a choice is completed, the orchestrator moves to a state where only an action or termination can be chosen. After an action is completed, the orchestrator returns to the Start location. The choice can be either dictatorial or majoritarian, according to the assigned configuration. In the majoritarian choice, the orchestrator interacts with the services to collect their choice. Similarly, the configuration of action can be either centralised or distributed. In both cases, an offer label or a match label will be executed. As stated in Section 3, in a distributed match the orchestrator makes the two services aware of each other so that they can interact and execute the match. In case of termination, the orchestrator sends the termination message to all services and terminates.
5.1.2. Detailed description
We now briefly describe Figure 9 and Figure 10. In the first transition of the orchestrator, one of four options encoding the combinations of choice and action is selected non-deterministically. The function initialize(conf) instantiates accordingly the choice and action global variables such that all have the same configuration, and also initialises the communication buffers.
From the location CheckCompatibility of the orchestrator a loop is executed to communicate with all services to check if all have the same configuration. The orchestrator sends the message ORC_CHECK and its configuration (action and choice). If a service has the same configuration an ACK is sent, or an ERROR message otherwise. In case of no errors, the orchestration starts. As stated above, from the location Start the orchestrator internally decides whether to perform a choice, an action or terminate.
@@ 5.1.2. Detailed description
In the case of termination, the orchestrator sends to all services an ORC_STOP message and the orchestration terminates.
Choice
In the case of a choice, firstly all services receive the message ORC_CHOICE. If the choice is dictatorial, the orchestrator decides autonomously, and no further interactions are necessary. Otherwise, in the case of a majoritarian choice, the services involved in the choice are selected non-deterministically, and a message ORC_CHOICE is sent to them. Concretely, the involved services are those who perform an action in one of the outgoing transitions from the current state of the orchestration. The services that are not involved will receive a SKIP message. The involved services then receive from the orchestrator the available choices (i.e., the forward star of the current state). These concrete choices are abstracted by the constant CHOICES. Each involved service now replies with its choice, abstracted as a message SERVICE_CHOICE. After all involved services have voted, the orchestrator will decide accordingly.
Action
@@ 6. Analysis
We now describe the analyses we performed on the model. The modelling activity led to some issues in both the implementation and the model, which were all fixed (we remark that an already existing application was modelled). Other formal checks on the model were performed to ensure the properties described in this section (e.g., absence of deadlocks, absence of orphan messages). We have used Uppaal version 4.1.26-1, released in February 2022. Note that a more recent version of Uppaal is available, which integrates the previously separate tool Uppaal Stratego. We opted to continue using the version we initially started working with, since the capabilities of Uppaal Stratego are not used in this paper.
@@ 6.1. Validation through modelling
On a side note, thorough testing is generally more time-consuming than designing a formal model similar to the one in this paper. This is a further benefit derived from the usage of formal methods. For example, the library implementing contract automata operations (CATLib[13] has been tested up to 100% coverage of all lines and branches, also using mutation testing [13]. In CATLib, the lines of code of the tests are more than three times those of its source code [13]. A similar effort for CARE is more demanding than the one needed for designing the models in Figures 9 and 10. Uppaal has been used to automatically generate tests from the formal model.
@@ 6.2. Formal Verification
Performances. Statistical model checking has been used to scale to larger systems, and the verification has been performed in a few seconds on a standard laptop. Conversely, exhaustive model checking necessitated more resources and has been limited to smaller parameter setup generating hundreds of millions of states (see below). In this case, the verification has been performed on a machine with Intel(R) Core(TM) i9-9900K CPU @ 3.60 GHz equipped with 32 GB of RAM. Uppaal has been configured for maximising the state space optimisation and reusing the generated state space. Logs of the experiments are available in [10]. The summaries of the analysis performed with statistical (resp., exhaustive) model checking are in Table 1 (resp., Table 2).
@@
Table 2. A summary of the properties analysed with exhaustive model checking that are satisfied by the model
@@ 6.3. Parameters Tuning
The verification process involves employing specific parameter setup within the model. This encompasses various aspects, such as setting the delays in reading and writing, setting socket timeout thresholds, adjusting buffer sizes, assigning probability weights, and determining the number of services involved (i.e., the instantiations of the RunnableOrchestratedContract template). For deriving the desired set-up of parameters, we employ statistical model checking. If not stated otherwise, the parameters $\alpha$ and $\varepsilon$ of the statistical model checker are set to $0.05$ and $0.005$, respectively (see Section 3).
@@
It is crucial to note that we do not employ statistical model checking to determine values (such as buffers size) for use in the concrete implementation. Instead, these quantified values are used solely within the model. For instance, in the actual implementation, the size of Java TCP/IP socket buffers is fixed (more below). Our objective is to employ parameter setup that ensure realistic modelling and improve the performances of the exhaustive model checking. Realistic modelling entails accurately representing the behaviour of the real system, by reducing the probability of filling the buffers, timeouts, or excessive communication delays. Failure to maintain these conditions could potentially invalidate the results of the formal verification. Improving performances, on the other hand, entails reducing the state space of the model.
Probabilities weights. The values of the probabilities weights pstop, pchoice, pnochoice, and paction (see Figure 9) can be tuned based on average values extracted from the orchestrations subject of the analysis. Indeed, as stated in Section 4, the underlying orchestration automaton is abstracted away. Note that, e.g., in location Start the probability of performing a choice is $\frac{\tt pchoice}{\tt pchoice+pnochoice}$ (pchoice is the probability weight of performing a choice and pnochoice is the probability weight of not performing a choice). For example, the orchestration in Figure 11 (right) can be modelled by tuning the probabilities to pstop=25, pchoice=1, pnochoice=0, and paction=75.
Figure 11. On the left side, the SocketTimeout template automaton (one such automaton is istantiated for each service). On the right side, the orchestration automaton taken from the composition service example in [15]
For readability, in the reminder of this section we will use in the formulae the abbreviations ror for the instantiation of the RunnableOrchestration template (i.e., the orchestrator), and ROC(i) for the i-th instantiation of the RunnableOrchestratedContract template (i.e., a service).
Buffer size. Next, we address the buffer size (denoted as variable queueSize). It is important to note that the buffers in the model are represented by global arrays utilized by the automata for enqueuing and dequeuing values. These global arrays serve as models of the actual buffers found in Java TCP/IP sockets, where the default size is typically 8 KB. Our objective is to prevent unnecessary growth in the model’s state space while ensuring a low probability of the buffers filling up, similar to the behaviour observed in the real application. The formula:
@@
Delays and timeout. Next, we consider the real-time behaviour of the model and focus on determining appropriate values for the rates write and read, which represent the delays in writing to and reading from a buffer, respectively. The average message delay in Java TCP/IP sockets is affected by multiple factors, such as network conditions and server load. Delays can be sampled using tools like tcpdump, and the rate can be estimated by calculating the inverse of the sample mean. Additionally, we consider the variable timeout, which represents the timeout threshold. Our aim is to achieve three objectives. First, we strive to maintain a low probability of encountering timeouts. Second, we seek to ensure a high probability of terminating within a specific timeframe, which we have set to be 500 time units, corresponding to the duration used in our experimental setup. Third, we aim to keep the timeout threshold at a lower value in order to decrease the state space and facilitate model checking.
@@
Concerning the instances of the templates, in all experiments there is one instance of the orchestrator template and either 4 or 5 instances of the service template. For the exhaustive model checking phase, we used two small parameter setup of (number of services, buffers size). The first is c1=(4,5), the second is c2=(5,3). In fact, the parameter setup (5,4) remained inconclusive in the experiments due to the need for generating billions of states and the inadequate memory capacity of the utilized machine. The verification of larger parameter setup necessitates either relying exclusively on statistical model checking or employing more powerful machines.
@@ 6.4. Verification
Once the model’s parameter setup is determined, our next step involves verifying additional formal properties.
@@
i.e., if the orchestrator starts to terminate then eventually all services and the orchestrator terminate. The formula p-->q is a shortcut for A[](p imply A<>q). Thus, this formula states that for all executions and for all states either the orchestrator is not in the location ror.Stop or all executions passing through that location will eventually lead to a state where all services and the orchestrator have terminated or a timeout failure is experienced. The formula holds in both parameter setup c1 and c2. The first (resp. second) parameter setup required roughly 3 (resp. 30) minutes, 1.5 (resp. 12) Giga of memory, and explored 52 (resp. 426) million states.
@@
The formula states that for all executions and for all states of the composed system, either there is always at least one enabled transition or either a timeout failure has been experienced or all services and the orchestrator are in the Terminated location. In the formula, not deadlock is a special predicate provided by Uppaal. As expected, this property is satisfied in both parameter setup. The first (resp. second) parameter setup required roughly 3 (resp. 40) minutes, 1.5 (resp. 13) Giga of memory, and explored 25 (resp. 189) million states. This also proves that in a correct configuration (of action and choice) the Error location is never reached, because this would result in a deadlock and the above property would not be satisfied. This property is also satisfied when the size of the buffers is $3$. However, if we further reduce the size, then a deadlock occurs. This is because from location CheckCompatibility the orchestrator requires to insert three messages in the buffer of the receiver in one step. By dividing these three send operations in three non-committed transitions it is possible to further reduce the buffer size.
@@
this property quantifies the probability of termination with at least one message remaining in any buffer. To verify whether all buffers are empty, we utilize the predicate allEmpty(). As expected, the probability is found to be nearly zero. We proceed with an exhaustive verification by employing the property:
@@
the above formula can be read as follows: in all states of all executions either all buffers are empty or someone has not terminated yet. The above property is valid in both parameter setup, as expected. The first (resp. second) parameter setup required roughly 2 (resp. 27) minutes, 1.5 (resp. 12.5) Giga of memory, and explored 25 (resp. 189) million states. It is also possible to verify that there is no dummy execution in which the services and the orchestrator never interact (and thus all buffers are trivially empty). This can be verified with the formula:
@@
that, as stated above, checks if there exists an execution where all states have empty buffers (excluding the dummy execution scenario in which the timeout occurs at the beginning). As expected, this property is not valid in the model, for both parameter setup. The first (resp. second) parameter setup required roughly 3 (resp. 24) seconds, 1.4 (resp. 10) Giga of memory, and explored 24 (resp. 28) states.
No interference. When discussing the distributed match action in Section 5, we stated that it is never the case that some service interferes in a match in which it is not involved. This guarantees that it is safe to use two one-position buffers for all communications between any two services involved in a match. To verify this, we perform statistical model checking of the following formula:
@@ 7. Testing
The abstract test cases generated by Uppaal have been transformed into concrete JUnit tests for the source code. Uppaal provides various methods for test generation (see Section 3). We utilized the generation from reachability queries (i.e., of the form E<>). In our models, these queries encode specific simulation traces that are relevant to the specific orchestration employed in the tests and are designed to cover all transitions of the model. The specific simulation traces also guide the generation of the concrete test cases from the abstract test cases. Uppaal provides functionalities for offline test generation, such as achieving transition coverage or generating traces that reach a specified final state (reachability queries). However, Uppaal does not primitively offer a mechanism within their query language to explicitly enforce a sequence of intermediate states or events that a test trace must follow. To achieve this, two global variables, namely steps[] and step, are introduced in the models to encode the desired orchestration in the query. Whenever a specific transition is executed, the variable steps[step] is updated with a value encoding the transition being executed, and the counter step is incremented. For example (see Figure 12), whenever the transition reaching state Terminated is executed, the value ORC_STOP is assigned to steps[step]. As another example, whenever a centralised match transition is executed by the orchestration, the value CENTRALISED_MATCH is assigned to steps[step], encoding the execution of the request action in the match (see Section 5). The length of the array steps[] is equal to the desired number of transitions that the trace must pass through, and the variable step is a pointer to the current position in the array. As a small example, let alice be an instatiation of the automaton in Figure 12. The query E<>(alice.steps[0]==ORC_CHECK) encodes a trace where the check operation is executed first. In Section 7.3, we will provide a complete example showing how these additional variables are used in a query.